OpenStudy (anonymous):

Circle the first term in the trinomial. multiply it by each term in the binomial, and place each result in the spaces provided. Repeat the proccess for the second and third term of the trinomial until all of the spaces are filled in. (3x^3-2x^2+7)(x^3+2x) ___+____+____+_____+____+____

OpenStudy (fibonaccichick666):

ok so remember how I said foil is a lot like the distributive property?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

so if it was just (3x^3-2x^2+7)(x^3) what would you do?

OpenStudy (anonymous):

ohhhh

OpenStudy (fibonaccichick666):

then if you just had (3x^3-2x^2+7)(2x) what would you do?

OpenStudy (fibonaccichick666):

add em together and what do ya get?

OpenStudy (anonymous):

whered you get (2x)

OpenStudy (fibonaccichick666):

\[(3x^3-2x^2+7)(x^3+\color\red{2x})\]

OpenStudy (fibonaccichick666):

I prefer to do less work so I rearranged a little

OpenStudy (fibonaccichick666):

so if you have a(b+d)=ab+ad right? same idea here

OpenStudy (anonymous):

i did it can you check if i am correct?

OpenStudy (fibonaccichick666):

yeppers

OpenStudy (fibonaccichick666):

don't skip any steps please

OpenStudy (anonymous):

3x^9+6x^3-2x^5-4x^2+7x^3+14x

OpenStudy (anonymous):

oh dammit

OpenStudy (anonymous):

hold on ama do the work

OpenStudy (fibonaccichick666):

also when you do this multiplication, what do you do with the exponents? \[x^3*x^3=?\]

OpenStudy (anonymous):

add them

OpenStudy (fibonaccichick666):

yea, check what you did up there before you type up your thing